Express the function g:R_R. Defined by th g(x) as the composition of the three
1
Expert's answer
2020-04-04T16:05:35-0400

"g(x)=\frac {1}{\sqrt{x+\sqrt{x}}}," "g:R^+\to R^+"

Let "f(x)=\sqrt{x}"

"h(x)=\sqrt{x^2+x}"

"k(x)=\frac 1 x"

We find:

"h(f(x))=\sqrt{x+\sqrt{x}}"

Again:

"k(h(f(x)))=\frac{1}{\sqrt{x+\sqrt{x}}}"

So we concluded that

"g(x)=k(h(f(x)))," where "f(x)=\sqrt{x};h(x)=\sqrt{x^2+x};k(x)=\frac 1 x"